From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

  • Location:
  • ESNET: 8242553
  • Phone:
    • +1-800-377-8846 : US
    • +1-888-276-7715 : Canada
    • +1-302-709-8424 : International
    • Enter Code: 39527048#
  • Time: 10:00am EST




  • JLab: Tim, Simon, Dave L., Carl, Eugene, Mark, John D., Dan, Elke, Claus, Dave D., Fernando
  • UConn: Richard
  • IUCF: Gerard
  • CMU: Curtis, Yves
  • Athens: Christine


Elke had a few:

  • JLab in a hiring freeze at present
  • Next week, Monday trought Friday, is the EVMS review
    • cost account managers will make presentations
    • physics is scheduled for Wednesday morning

Elke proposed that we drop our meeting for next week. All agreed.

R&D Contracts

Elke reported on this item.

  • Work proceeds in conjuction with JLab Procurement
  • Several contracts that ran from 2002 to 2007 have to be re-visited
  • These include IUCF and Regina - The Regina one has been dealt with: second last invoice has been sent, last one will be after Phase-1 close out before the end of the year
  • Most likely these contracts will be closed and re-open with new ones.
  • Elke has list of all known contracts
    • It is an Excel spreadsheet
    • She will send it around soon.

Report from Calorimetry Technical Group

Zisis reported on items from the last meeting of the group. See his minutes for details.

Later today, Tim will send preliminary drawings of the BCAL wedges and assembly to George and Zisis, so they can see whether the Alberta/Regina machine shops can constructed these.

Detector Updates

Beamline and Tagger

  • Elke: design contract for tagger to go to procurement soon
    • will be reviewed by accelerator division
  • Elke: documents on the wiki are near up-to-date
  • Tim: sample for tagger interface drawing will be sent this week


Elke reported items from George B.:

  • coil 3 looks worse that we thought
    • will completely open and replace some tubing
    • there are some contract issues with IUCF to be sorted out (see above)
  • we have started testing coils at JLab


Tim reported on yesterday's meeting.

  • reviewed tracking list
  • temporary solid target discussed (for use during startup)
  • specification to-do list reviewed
  • discussion of racks for the gas system, gas system itself, were discussed
  • discussion of alignment accuracies
  • discussion of thermometry
  • discussion cryogenics: independent system or use magnet system

Start Counter

Werner reported on discussions of geometry. Some items:

  • overall dimensions, stay-in zone
  • target independent of start counter desired
  • 2004 geometry for counters themselves still looks good with current rate estimates


Yves gave a detailed report on recent progress at CMU. See his status reports (1, 2) on the web for details.

Gerard had some concerns about cabling and grounding based on the photograph of the prototype. He will talk with Yves offline.

Gerard's trip to CMU has yet to be scheduled. He wants to look at some of the CMU electronics in advance of the trip.


Dan gave the report:

  • Procurement of the wire winding contract is waiting on final budget from IUCF
  • Composite wire frames are being produced. Brian will have five by next week.
    • construction is more accurate with new jig
  • Test Lab set-up being put back together. Denny Insley is helping.
  • another offer of new space for FDC's downstairs in CEBAF Center has materialized
    • Brian also has space in the EEL.
    • Space in the clean room being discussed with Jack Segal
  • Interposer/daughter boards have not yet been transferred from Fernando. FDC group not quite ready to test them yet.
  • There will be an FDC meeting tomorrow; agenda has been circulated
  • The combined FDC/CDC/tracking meeting TBA


Zisis gave the report:

  • Tests with cosmics with SiPM's on the BCAL continue.
  • Monte Carlo work continuing.
  • George L. is working on the statement of work for the SiPM's
  • Elke and Carl: test fibers have been ordered from St. Gobain (Bicron). The order from Kuraray should go out this week.
    • Carl: it might be two months to get them without some special effort or arrangement. Elke encouraged Carl to see if this can be sped up, since otherwise it would be outside our BCAL readout timelines.


Beni gave the report:

  • The test set-up for the plastic wall for holding PMT's has been constructed.
  • There will be a CD3 preparation meeting Friday at 8:30 am.
  • He will be getting NIM electronics from an IUCF transportainer. He hopes they work.
  • Zisis: George L. has asked about beamlines at TRIUMF for tests (electrons from ~80-150 MeV for FCAL threshold test)
    • beamline M11 (medium energy pion beam, with muons and electrons from decays) is available
    • need two month notice to get time on it
    • the 8x8 block FCAL test is a candidate
    • electronics/read-out system must come with the experimenters; little or no electronics are available from the TRIUMF pool
    • these tests will use the JLab online DAQ system
  • Beni has made inquiries about Russian rad-hard lead glass blocks
    • 1 block for $260
    • limited accuracy in dimensions: 2 mm tolerance


Christina reported:

  • They are collecting components for a prototype system.
  • They have initiated discussions with a company for purchasing LED drivers.
  • They have some tiny LED's they will use to begin work.


Fernando reported:

  • Pre-amp cards, cables and connectors for the FDC and CDC will be done this week.
    • They are assembled; testing remains to be done.
  • The 250 MHz FADC is in good shape.

Monte Carlo

Richard reported on the new geometry. See his wiki page for details. He remarked in several items:

  • The space between stereo and axial packages in both versions may be able to be reduced.
  • The "endplate" material of the CDC is actually dominiated by the endplugs for the straws. These plugs need to be modeled properly; this is not the case presently.
  • He requested feedback on BCAL geometry details.